A ball is thrown into the air vertically from the ground and falls back to the
ground 2.0 s later. The acceleration due to gravity is 9.8 m/s2. How high did the ball reach?
1
Expert's answer
2021-10-12T12:32:24-0400

The maximum height can be found with the following equation:


"h=\frac{g(T/2)^2}{2}=\frac{9.8(2/2)^2}{2}=
4.9\text{ m}."